OpenStudy (anonymous):

last Question for the night guys will medal, 4√16x^4

Nnesha (nnesha):

did you understand last question ??

OpenStudy (anonymous):

Not really to be honest.

Nnesha (nnesha):

so you have to learn the method instead of taking just answer this one is totally same like last one same process with different number

Nnesha (nnesha):

if you look at ur previous post i hope you will be able to do this one try it :)

OpenStudy (anonymous):

thank you so much I will(:

Nnesha (nnesha):

let me know what you get ???

OpenStudy (anonymous):

I will

Nnesha (nnesha):

after solving this i'll wait :)

OpenStudy (anonymous):

I dont understand it @Nnesha I tried

Nnesha (nnesha):

okay did you start doing anything with this problem ??? how fare u get that

Nnesha (nnesha):

can you chang 4th root to exponent od (16x^4) like we i did on previous post

OpenStudy (anonymous):

so i followed the steps and rewrote the way you rewrite the other one

Nnesha (nnesha):

yes same like last one

Nnesha (nnesha):

yes so what you got so fr

OpenStudy (anonymous):

(16x^4)^1/4?

Nnesha (nnesha):

yes that's right now you can separate 16 and x^4\[\huge\ (16)^{\frac{ 1 }{ 4 }} \times (x^4)^{\frac{ 1 }{ 4 }}\]

OpenStudy (anonymous):

ok

Nnesha (nnesha):

so what is \[\huge\ (16)^{\frac{ 1 }{ 4 }}=???\]

OpenStudy (anonymous):

4?

Nnesha (nnesha):

nope what is 2 times 2 times 2 times 2

OpenStudy (anonymous):

16

Nnesha (nnesha):

yes so you can write that as 2^4 which is = 16 so replace 16 by 2^4 now what is

Nnesha (nnesha):

\[\huge\ (2^4)^{\frac{ 1 }{ 4 }}\]

Nnesha (nnesha):

^^^^ = ???

OpenStudy (anonymous):

I have no idea

Nnesha (nnesha):

 multiply inside exponent by outside exponent